# Wenfan Ou and Sheng Bi

### v3

@@ A $\phi$-divergence ambiguity set is defined via
Also utilizing $\chi^{2}$-divergence, Klabjan et al. (2013) directly use historical data to construct ambiguity sets in a goodness-of-fit test paradigm. Let the number of observations falling within the i-th bin be a function of a realized random variable $\xi_{t}\sim\mu_{t}$ as $N_{t,i}(\xi_{t})$, and the total number $n_{t}(\xi_{t})=\sum_{i}N_{t,i}(\xi_{t})$. The ambiguity set consists of the distributions $\mu_{t}$ which satisfy
$\sum_{i}\frac{(N_{t,i}(\xi_{t})-n_{t}(\xi_{t})\cdot\mu_{t}(i))^{2}}{n_{t}(\xi_{t})\mu_{t}(i)}\leq\theta,\quad t=1,...,T,$ (22)
where $\hat{\mu}_{t}(i)$ denotes the estimated probability that the number of observations falling within the $i$-th bin, and parameter $\theta$ controls how close the observed sample data is to the estimated expected number of observations according to the fitted distribution $\mu_{t}$.
Notably, the literature on RMDPs with $\phi$-divergence ambiguity sets has recently paid more attention to sample complexity and efficient algorithm design in RL. In these aspects, there are also interesting and impressive results. We briefly review several advanced literature in the following.
For sample complexity analysis, Wang, Si, Blanchet & Zhou (2023) extend Liu et al. (2022), and first provide sample complexity result for the model-free robust RL problem. Shi et al. (2023) investigate RMDPs with the uncertainty set measured via total variation distance. Under mild conditions, they prove that RMDPs in this case are more sample-efficient than standard MDPs via model-based algorithms. For efficient algorithms design, Grand-Clément & Kroer (2021) attempt to replace the original value function with the first order approximation in robust Bellman update, while Ho et al. (2022) employ bisection method to solve robust Bellman update for commonly $\phi$-divergence, e.g., KL-divergence, Burg entropy, Variation distance, and $\chi^{2}$ distance.
@@ 5.3 Wasserstein Distance Ambiguity
Although $\phi$-divergence has indeed performed well in some cases, there are some obvious limitations to obstacle the flexibility of modeling with $\phi$-divergence, e.g. only working on the same support and not satisfying the triangle inequality. Nowadays, Wasserstein distance is deemed as a good substitute for $\phi$-divergence. Wasserstein distance is a particular case of optimal transport discrepancies, which computes the minimal cost of transporting the masses between two distributions. Based on the definition of Wasserstein distance, the Wasserstein distance ambiguity is defined in the following.
@@
A Wasserstein ambiguity sets can be defined via
$\mathcal{P}\triangleq\left\{P\in\mathfrak{M}(\Xi,\mathcal{F})\left|~{}\mathcal{W}(P,\hat{P})\leq\theta\right\}\right..$ (23)
Wasserstein distance $\mathcal{W}(P,\hat{P}):=\min\limits_{\Gamma\in\mathfrak{M}\times\mathfrak{M}}\left\{\int_{\Xi\times\Xi}c(\xi,\zeta)\Gamma(\mathrm{d}\xi,\mathrm{d}\zeta):~{}\Pi^{1}_{\#}\Gamma=P,\Pi^{2}_{\#}\Gamma=\hat{P}\right\}$, where $c(\xi,\zeta)$ is the cost function of moving the mass from $\xi$ to $\zeta$, $\Pi^{i}_{\#}\Gamma$ denote the $i$-th marginal distribution of $\Gamma$, and $\Gamma(\cdot,\cdot)$ is the joint distribution of $P$ and $\hat{P}$, also called transport plan.
In addition to its representation in joint distribution form, Wasserstein distance also has a dual representation, which allows for flexibility in modeling in the context of RMDPs.
@@ A Wasserstein ambiguity sets can be defined via
Wasserstein distance has a dual representation due to Kantorovich duality (Villani 2009), and we can represent Wasserstein distance equivalently as
$W(P,\hat{P})=\sup_{\begin{subarray}{c}\psi\in L^{1}(P)\\ \varphi\in L^{1}(\hat{P})\end{subarray}}\left\{\int_{\Xi}\psi(\xi)P(\mathrm{d}\xi)+\int_{\Xi}\varphi(\zeta)\hat{P}(\mathrm{d}\zeta):~{}\psi(\xi)+\varphi(\zeta)\leq c(\xi,\zeta)~{}\forall\xi,\zeta\in\Xi\right\},$ (24)
where $L^{1}(P)$ denotes the $L^{1}$ space of $P$-measurable functions.
Prominent results from recent Wasserstein DRO literature (Esfahani & Kuhn 2018, Blanchet & Murthy 2019, Gao & Kleywegt 2023) provide normative reformulations and finite-sample guarantees, facilitating to construct ambiguity sets in a data-driven manner and inspire the research on RMDPs with Wasserstein ambiguity sets.
Yang (2017) first investigates the Wasserstein RMDPs. The author constructs the Wasserstein ambiguity sets over each state $s\in\mathcal{S}$, falling within $s$-rectangularity implicitly, and derives corresponding distributionally robust Bellman equations by applying the dynamic programming principle,
@@
where $(P_{s},R_{s})$ obeys to a joint distribution $\mu$. The author then derives an equivalent and tractable reformulation by leveraging the strong duality of the proposed Bellman equation.
As the extension of Yang (2017), Yang (2020) considers a continuous-state setting. Besides providing the tractable reformulation with no duality gap and iteration algorithms, the author tailors more specific properties from Wasserstein DRO literature to characterize proposed Wasserstein RMDPs. Under some mild assumptions, the optimal policy of the worst-case is deterministic and stationary, and its structure can be expressed explicitly as in Gao & Kleywegt (2023). Moreover, a probabilistic out-of-sample performance guarantee with mild assumptions is established as in Esfahani & Kuhn (2018).
Subsequently, Ramani & Ghate (2022) consider RMDPs with general distance metrics, such as total variation (TV), Burg, Hellinger, and Wasserstein distance. The authors focus on the ambiguity sets where any state-action-stage triplet is a ball, with $(s,a)$-rectangular assumption to maintain generality. They prove the optimal values of the RMDPs converge almost surely to the true optimal value as the sample size diverges to infinity, and also provide finite-sample probabilistic performance guarantee on out-of-sample values.
Besides resorting to general dynamic programming methods, there is increasing literature combining Wasserstein distance with RL algorithms. Different from previous works that merely employ Wasserstein ambiguity sets to obtain robust actions (Kandel & Moura 2020), Neufeld & Sester (2022) design a robust Q-learning algorithm to directly construct ambiguity sets, established above $(s,a)$-rectangular assumption implicitly. By introducing $\lambda c$-transform, which is proved to be equivalent to the duality of the value objective function, the authors reformulate the Q-value update as a regularized form.
Hou et al. (2020) design a Wasserstein Robust Advantage Actor-Critic algorithm (WRAAC), and leverage the strong duality proof (Blanchet & Murthy 2019) to convert the original problem into a finite-dimensional ordinary risk-aware RL problem. Through the moderated optimal Bellman equation, they prove the existence of robust optimal policies and provide the theoretical analysis of the performance of optimal policies. Based on the above theoretical foundations, the authors develop the WRAAC algorithm via two neural networks to estimate value function and approximate the action policy. WRAAC is more practical than the Wasserstein robust value iteration proposed by Yang (2017), as it can approximate nominal distribution at each step without estimating enough transition kernels.
6 Beyond Rectangularity: Coupled Uncertainty Modelling

### v4

@@ A $\phi$-divergence ambiguity set is defined via
Also adhering to $\chi^{2}$-divergence, Klabjan et al. (2013) directly utilize historical data to construct ambiguity sets from a goodness-of-fit test perspective for the single-item multi-period periodic review stochastic lot-sizing problem. Different from the approach in Hanasusanto & Kuhn (2013), the authors construct the nominal distribution discretely. Let $B$ represent the set of bins that partition the range of possible values of the random variable $\xi_{t}\sim\mu_{t}$. For each $i\in B$, let $N_{t,i}(\xi_{t})$ be the number of observations falling within the $i$-th bin, and the total number $n_{t}(\xi_{t})=\sum_{i\in B}N_{t,i}(\xi_{t})$. The ambiguity set consists of the distributions $\mu_{t}$ which satisfy
$\sum_{i\in B}\frac{(N_{t,i}(\xi_{t})-n_{t}(\xi_{t})\cdot\hat{\mu}_{t}(i))^{2}}{n_{t}(\xi_{t})\hat{\mu}_{t}(i)}\leq\theta,\quad t=1,...,T,$ (34)
where $\hat{\mu}_{t}(i)$ denotes the estimated probability that an observation falls within the $i$-th bin. The parameter $\theta$ controls how close the observed sample data is to the estimated expected number of observations according to the fitted distribution $\mu_{t}$. From a hypothesis testing perspective, the authors set $\theta=\chi^{2}_{|B|-1,1-\alpha}$ where $|B|$ represents the total number of bins and $\alpha$ denotes the significance level. Consequently, all distributions satisfying the constraints in (34) are those for which the corresponding null hypothesis is not rejected at the significance level $\alpha$. By implementing a standard dualization process, the original robust problem can be converted into a tractable second-order cone program. In this paper, the authors show that a state-dependent $(s,S)$ policy is optimal for this robust lot-sizing problem and the ordering levels can be computed by this second-order cone program.
Notably, recent literature on RMDPs with $\phi$-divergence ambiguity sets has increasingly focused on sample complexity and the design of efficient algorithms. Unlike the efficient algorithms for RMDPs with norm ambiguity sets, which typically necessitate the use of LPs, the efficient algorithms involving $\phi$-divergence display distinct methodologies. In these aspects, there have been several noteworthy and impressive advancements. Below, we briefly review some of the advanced literature in this area.
For sample complexity analysis, Wang et al. (2023) extend the distributionally robust Q-learning framework proposed by Liu et al. (2022) through refining the design and analysis of the key component, multi-level Monte Carlo estimator. Specifically, while the expected number of samples requested in Liu et al. (2022) is infinite and the algorithm only can achieve asymptotic guarantees, the refined algorithm in Wang et al. (2023) requires a merely constant order number of samples. This improvement is achieved by delicately devising the sample numbers for each state-action pair. This is also the first sample complexity result for the model-free robust RL problem. Shi et al. (2023) investigate RMDPs with the uncertainty set measured via total variation (TV) distance or $\chi^{2}$-divergence. The authors suppose that one has access to a generative model or simulator to draw samples with a nominal transition probability matrix. Under $(s,a)$-rectangular assumption, they propose a model-based distributionally robust value iteration algorithm. In each iteration, the robust Q-function is computed in terms of each state-action pair, and the robust value function is set greedily according to the current robust Q-function. Interestingly, the authors find that the choice of uncertainty set significantly impacts the sample size requirements. For instance, to achieve the same $\epsilon$-accuracy, while the RMDPs with $\chi^{2}$-divergence are harder than standard MDPs as expected, the RMDPs with TV distance are easier than standard MDPs. These findings emphasize the importance of the construction of uncertainty sets and the efficiency of the model-based approaches.
For efficient algorithms design, Grand-Clément & Kroer (2021) introduce the first first-order method (FOM) for solving RMDPs and propose a scalable algorithmic framework of FOM updates with occasional approximate value iteration updates (FOM-VI) to limit the computational cost of value iteration. This framework represents the first tractable algorithm for $s$-rectangular RMDPs with KL-divergence ambiguity sets. It is noteworthy that previously, only the $(s,a)$-rectangular case had been effectively addressed by robust value iteration in Nilim & El Ghaoui (2005). A key observation underpinning this algorithm is that an $s$-rectangular RMDP can be decomposed into $|\mathcal{S}|$ bilinear saddle-points problems (BSPPs). Within each epoch, a primal-dual algorithm (PDA) is employed to update the action policy $\pi_{s}$ (primal updates) and the conditional transition matrix $\mathbf{p}_{s}$ (dual updates) with fixed value-to-go function $V_{s}$ for each state $s$ (i.e., a BSPP). At the end of this epoch, $V_{s}$ is updated by fixed policy and conditional transition matrix, which are constructed by a weighted average of the policies or conditional transition matrices with designed weights. A novel feature of this algorithm lies in the PDA updates, where the associated proximal mappings are defined as follows:
$\mathrm{prox}_{\pi}(\bm{g}_{\pi},\pi_{s}^{\prime})$ $=\arg\min_{\pi_{s}\in\Delta^{|\mathcal{A}|}}\langle\bm{g}_{\pi},\pi\rangle+D_{\pi}(\pi_{s},\pi_{s}^{\prime}),$ (35)
$\mathrm{prox}_{\mathbf{p}}(\bm{g}_{\mathbf{p}},\mathbf{p}_{s}^{\prime})$ $=\arg\max_{\mathbf{p}_{s}\in\mathcal{P}_{s}}\langle\bm{g}_{\mathbf{p}},\mathbf{p}_{s}\rangle-D_{\mathbf{p}}(\mathbf{p}_{s},\mathbf{p}_{s}^{\prime}),$
where $\bm{g}_{\pi}$ and $\bm{g}_{\mathbf{p}}$ are the gradients with respect to current $\pi_{s}$ and $\mathbf{p}_{s}$, respectively, and $D_{\pi}(\cdot,\cdot)$ and $D_{\mathbf{p}}(\cdot,\cdot)$ are Bregman divergence functions. The intuition behind (35) is that proximal mappings move along with the direction of improvement as indicated by the gradients while being penalized by the Bregman divergences. These penalties ensure that the updates remain within a region where the first-order approximations are accurate enough. Given the KL-divergence constraints in $\mathcal{P}_{s}$, the authors show that there exists a Bregman divergence such that solving the original RMDPs is equivalent to operating the FOM-VI framework, where the PDA updates are governed by the proximal mappings in (35).
Ho et al. (2022) further generalize existing literature by proposing a fast algorithmic framework to solve general $\phi$-divergence RMDPs for both $(s,a)$-rectangular and $s$-rectangular ambiguity sets. Unlike Grand-Clément & Kroer (2021), which replaces traditional VI procedures with FOMs updates, the authors adhere to previous research methodologies in RMDPs, focusing on accelerating the robust Bellman update in robust VI. The key challenge in solving $s$-rectangular RMDPs with $\phi$-divergence ambiguity sets lies in the fact that the corresponding ambiguity sets are non-polyhedral. The non-polyhedral nature makes each iteration (i.e., the robust Bellman update) computationally costly, a problem not present in the $(s,a)$-rectangular counterparts. This explains why the $(s,a)$-rectangular case can be efficiently solved by robust VI, whereas the $s$-rectangular case cannot. Given this barrier, a crucial component in this paper that circumvents the aforementioned challenge is the authors’ discovery that these min-max problems can be reduced to a small number of highly structured projection problems onto a probability simplex.
Specifically, by applying the minimax theorem, the optimal value of the original max-min problem equals to min-max problem where the inner problem is optimized over a deterministic worst action, rather than a (randomized) policy. The min-max problem can be efficiently solved via the bisection method; which finds the lowest possible constant $\beta$ such that the objective value is not larger than $\beta$ for any feasible solution. Once $\beta$ is given by the bisection method, the following generalized $D_{\phi}$-projection problem of nominal transition probabilities $\hat{\mathbf{p}}_{sa}$ will be checked:
$\begin{array}[]{llll}\mathrm{P}(\hat{\mathbf{p}}_{sa};\bm{b},\beta)=&\mathrm{min}&&D_{\phi}(\mathbf{p}_{sa},\hat{\mathbf{p}}_{sa})\\ &\mathrm{s.t.}&&\bm{b}^{\top}\mathbf{p}_{sa}\leq\beta\\ &&&\mathbf{p}_{sa}\in\Delta^{|\mathcal{S}|}\end{array}$ (36)
where $D_{\phi}$ is the predetermined $\phi$-divergence function and $\bm{b}\in\mathbb{R}^{|\mathcal{S}|}_{+}$ is a known parameter vector. Set $\bm{b}=\bm{r}_{sa}+\lambda\bm{v}$ and compute the value $\sum_{a\in\mathcal{A}}\mathrm{P}(\hat{\mathbf{p}}_{sa};\bm{r}_{sa}+\lambda\bm{v},\beta)$ and the corresponding optimal solution $\mathbf{p}^{*}_{sa}$. Then, the authors distinguish between the following two cases:
  1. 1.
    If $\sum_{a\in\mathcal{A}}\mathrm{P}(\hat{\mathbf{p}}_{sa};\bm{r}_{sa}+\lambda\bm{v},\beta)\leq\theta$, then $\mathbf{p}_{s}=(\mathbf{p}^{*}_{sa})_{a\in\mathcal{A}}$ is a feasible solution to (36). Consequently, $\beta$ upper bounds the optimal objective value of the original RMDP.
  2. 2.
    If $\sum_{a\in\mathcal{A}}\mathrm{P}(\hat{\mathbf{p}}_{sa};\bm{r}_{sa}+\lambda\bm{v},\beta)>\theta$, then there is no feasible $\mathbf{p}_{s}\in\Delta^{|\mathcal{S}|\times|\mathcal{A}|}$ such that the objective value attains $\beta$ or less. Thus, $\beta$ becomes a lower bound of the optimal objective value.
As $\mathrm{P}(\hat{\mathbf{p}}_{sa};\bm{r}_{sa}+\lambda\bm{v},\beta)$ represents the lowest deviation between $\hat{\mathbf{p}}_{sa}$ and $\mathbf{p}_{sa}$ that satisfies the value-go-function is bounded by $\beta$ for each state-action pair $(s,a)$, the $\sum_{a\in\mathcal{A}}\mathrm{P}(\hat{\mathbf{p}}_{sa};\bm{r}_{sa}+\lambda\bm{v},\beta)$ is the total deviation between $\hat{\mathbf{p}}_{s}$ and $\mathbf{p}_{s}$. If case 1 holds, the optimal solution $\mathbf{p}_{s}$ is a feasible solution in which each $\mathbf{p}_{sa}$ is the worst-case. By duality theory, $\beta$ becomes the upper bound of the optimal value. In turn, if the deviation $\sum_{a\in\mathcal{A}}\mathrm{P}(\hat{\mathbf{p}}_{sa};\bm{r}_{sa}+\lambda\bm{v},\beta)$ exceeds the parameter $\theta$, the solution $\mathbf{p}_{s}$ is out of the constructed ambiguity set, and corresponding $\beta$ provides a lower bound. Given this analysis and relationship, consequently, if the projection problem like (36) can be solved efficiently, the robust Bellman update of RMDPs can also be computed efficiently by the bisection method.
@@ 5.3 Wasserstein Distance Ambiguity
While $\phi$-divergence has demonstrated strong performance in certain applications, it also has notable limitations that hinder its modeling flexibility. For instance, $\phi$-divergence operates only on distributions with the same support and does not satisfy the triangle inequality. In contrast, Wasserstein distance has emerged as a good substitute for $\phi$-divergence. Wasserstein distance is a particular case of optimal transport discrepancies, which computes the minimal cost of transporting the masses between two distributions. Based on the definition of Wasserstein distance, the Wasserstein distance ambiguity is defined as follows.
@@
The Wasserstein distance $\mathcal{W}(\mathbb{P},\hat{\mathbb{P}})$ of ,^$\mathbb{P},\hat{\mathbb{P}}\in\mathfrak{M}(\Xi,\mathcal{F})$ is defined via
$\mathcal{W}(\mathbb{P},\hat{\mathbb{P}})=\min\limits_{\Gamma\in\mathfrak{M}\times\mathfrak{M}}\left\{\int_{\Xi\times\Xi}c(\xi,\zeta)\Gamma(\mathrm{d}\xi,\mathrm{d}\zeta):~{}\Pi^{1}_{\#}\Gamma=\mathbb{P},~{}\Pi^{2}_{\#}\Gamma=\hat{\mathbb{P}}\right\},$ (37)
where $c(\xi,\zeta)$ is the symmetric cost function of moving the mass from $\xi$ to $\zeta$ (i.e., $c(\xi,\zeta)=c(\zeta,\xi)$), $\Pi^{i}_{\#}\Gamma$ denote the $i$-th marginal distribution of $\Gamma$, and $\Gamma(\cdot,\cdot)$ is the joint distribution of $\mathbb{P}$ and $\hat{\mathbb{P}}$, also called the transport plan. Consequently, given the radius $\theta$ that controls the level of robustness, a Wasserstein ambiguity set can be constructed as
$\mathcal{P}\triangleq\left\{\mathbb{P}\in\mathfrak{M}(\Xi,\mathcal{F})\ \middle|\ \mathcal{W}(\mathbb{P},\hat{\mathbb{P}})\leq\theta\right\}.$ (38)
While the Wasserstein distance can be defined well in terms of a joint distribution $\Gamma$ in (37), its dual form, also referred to as the Kantorovich metric, sometimes contributes to the modeling and analysis. Formally, we claim that the following result is equivalent to the primal definition of Wasserstein distance in (37).
@@
Wasserstein distance has a dual representation due to Kantorovich duality (Villani 2009), and we can represent Wasserstein distance in (37) equivalently as
$W(\mathbb{P},\hat{\mathbb{P}})=\sup_{\begin{subarray}{c}\psi\in L^{1}(\mathbb{P})\\ \varphi\in L^{1}(\hat{\mathbb{P}})\end{subarray}}\left\{\int_{\Xi}\psi(\xi)\mathbb{P}(\mathrm{d}\xi)+\int_{\Xi}\varphi(\zeta)\hat{\mathbb{P}}(\mathrm{d}\zeta):~{}\psi(\xi)+\varphi(\zeta)\leq c(\xi,\zeta)~{}\forall\xi,\zeta\in\Xi\right\},$ (39)
where $L^{1}(\mathbb{P})$ denotes the $L^{1}$ space of functions that are $\mathbb{P}$-measurable. Especially, when the $c(\cdot,\cdot)$ is 1-order, namely, 1-Wasserstein distance is considered, the dual form can be further simplified as
$W(\mathbb{P},\hat{\mathbb{P}})=\sup_{\psi\in L^{1}(\mathbb{P})}\left\{\int_{\Xi}\psi(\xi)\mathbb{P}(\mathrm{d}\xi)-\int_{\Xi}\psi(\xi)\hat{\mathbb{P}}(\mathrm{d}\zeta)\right\},$ (40)
The dual formulation transforms the original minimization problem over transport plans $\Gamma$ into a maximization problem over potential functions $\psi,\varphi$, subject to a point-wise constraint. The dual form simplifies the original complex problem with coupling constraints as an optimization over 1-Lipschitz functions. Additionally, the dual formulation provides a framework for establishing theoretical results, such as convergence rates.
Prominent results from recent Wasserstein DRO literature (Esfahani & Kuhn 2018, Blanchet & Murthy 2019, Gao & Kleywegt 2023) offer tractable reformulation procedures and finite-sample guarantees, demonstrating that how to construct Wasserstein ambiguity sets in a data-driven manner and inspiring further research on RMDPs with Wasserstein ambiguity sets.
The Wasserstein DRMDPs were first investigated by Yang (2017), where the elements of the ambiguity sets are the joint distribution of transition probabilities and immediate rewards. Implicitly assuming $s$-rectangularity, the author derives the corresponding distributionally robust Bellman equation by applying DP principles
@@
and shows that this DRMDP admits an optimal Markov policy, while its construction is computationally challenging. To address the computational issue, the author rewrites the Wasserstein distance as the dual form and leverages Kantorovich duality (Villani 2021) to obtain a dual reformulation. Using this reformulation, the finite horizon DRMDPs can be solved by finite-dimensional convex programming. Motivated by the findings in Gao & Kleywegt (2023), the author offers a closed form of the worst-case probability distribution when the nominal distribution has finite support. However, no efficient algorithm or supplementary statistical properties, such as finite-sample guarantees, associated with Wasserstein distance are provided in this paper.
Building on the foundation work of Yang (2017), Yang (2020) considers a distributionally robust optimal control problem and extends the finite state space setting of Yang (2017) into a continuous state space setting. Although optimal control problems slightly differ from MDPs, as previously mentioned, they can be connected under some mild settings. Consequently, employing similar procedures and methodologies, the author also shows the existence of an optimal policy and provides a tractable dual reformulation with no duality gap. Given the $s$-rectangular assumption, the author employs value iteration and policy iteration algorithms within this dual problem framework, both of which involve solving semi-infinite programs in each iteration. The author also establishes the least number of iterations required to attain an $\epsilon$-optimal policy. Meanwhile, the author tailors specific properties from Wasserstein DRO literature to characterize proposed Wasserstein RMDPs. Under certain mild assumptions, the optimal policy for the worst-case scenario is deterministic and stationary, and its structure can be expressed explicitly, as shown in Gao & Kleywegt (2023), where Kantorovich duality and DP play a critical role. Furthermore, a probabilistic out-of-sample performance guarantee with mild assumptions is established, similar to the results in Esfahani & Kuhn (2018).
Apart from DRMDPs with Wasserstein distance, Ramani & Ghate (2022) consider $(s,a)$-rectangular RMDPs with general distance metrics under both finite horizon and infinite horizon settings, thereby including Wasserstein distance, where the nominal transition probabilities are derived by sample average approximation (SAA). As a theoretical paper, Ramani & Ghate (2022) aims to establish the following results: ($i$) robust value convergence, ($ii$) probabilistic performance guarantees on out-of-sample values, and ($iii$) a probabilistic convergence rate in the infinite horizon of rectangular RMDPs. For the first claim, the authors prove that the optimal values of the RMDPs converge almost surely to the true optimal value as the sample size $N\to\infty$. The key assumption required for the proof is the two probability mass functions are close if the metric function used in ambiguity sets deems them to be close. The intuition behind this assumption is that the simultaneous convergence of the empirical estimates to the true transition probabilities, along with the shrinking radius of the ambiguity balls, ensures convergence. Regarding the second claim, the rectangularity allows the probability of arbitrary guaranteed performance to be decomposed into the Cartesian product of the probabilities for each $(s,a)$ pair, facilitating further contractions. In the infinite horizon, by leveraging the simulation lemma in the RL literature (Rajeswaran et al. 2020), the authors bound the difference between the values of two policies through the corresponding transition matrices and thereby derive the third claim.
The successful application of Wasserstein distance in deep learning and classical RMDPs has inspired a growing body of research that integrates it with reinforcement learning algorithms. However, a considerable portion of literature adopts a two-step framework that treats Wasserstein distance as an independent plug-in component firstly, such as for updating simulator (Abdullah et al. 2019) or restricting feasible action set (Kandel & Moura 2020), and then solves a modified non-robust MDP. Not surprisingly, this two-step framework lacks theoretical guarantees and statistical properties that should be derived from the Wasserstein distance.
With stage-wise independent assumption, implying the $(s,a)$-rectangularity holds, Neufeld & Sester (2024) design a robust Q-learning algorithm with Wasserstein ambiguity sets. Different from previous works that utilize Wasserstein distance separately, the authors leverage the dual results of the robust Bellman equation due to $(s,a)$-rectangularity, and reformulate the Q-value update in a regularized form. Specifically, they introduce a so-called $\lambda c$-transform of function $f$, which is essentially a simplified and straightforward outcome resulting from strong duality as
$(f)^{\lambda c}(x):=\sup_{y\in\mathcal{X}}\{f(y)-\lambda\cdot c(x,y)\}$
where $\mathcal{X}$ is the support of $x,y$, $c:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is a predetermined cost function, and $f:\mathcal{X}\to\mathbb{R}$. In this paper, if we let $f$ be the value-to-go function and $c$ be the Wasserstein distance, the $\lambda c$-transform is the dualization procedure, omitting the Lagrangian multiplier term. This proposed algorithm is primarily rooted in the classical Q-learning algorithm. Nonetheless, before updating Q-values in each iteration, the optimal Lagrangian multiplier $\lambda^{*}$ will be computed such that the $\lambda c$-transformed Bellman equation with the Lagrangian multiplier term attains the supremum taken over the nominal probabilities:
$\lambda^{*}=\arg\sup_{\lambda\geq 0}\mathbb{E}_{\hat{\mathbf{p}}_{sa}}\left[-(-f_{(s,a)})^{\lambda c}(s^{\prime})-\varepsilon\lambda\right],$ (42)
where $f_{(s,a)}$ denotes the value-to-go function with respect to $(s,a)$. Subsequently, $\lambda^{*}$ will be used for the Q-values update where the value of dual form safely replaces the original (worst-case) Q-value, due to strong duality. The authors rigorously prove that the output of the proposed algorithm converges to the optimal robust Q-value function. Furthermore, the difference between the values of the robust and non-robust MDPs can be bounded if the Wasserstein ambiguity sets contain the true probability.
More generally, most RL settings lack a clear characterization of the environment, specifically in terms of access to a nominal distribution $\hat{\mathbf{p}}$. Consequently, model-based approaches, such as the one introduced by Neufeld & Sester (2024), become inapplicable (see the optimization procedure in (42) for reference). To address this issue, Hou et al. (2020) introduce a model-free algorithm called Wasserstein Robust Advantage Actor-Critic (WRAAC), which embeds Wasserstein constraints into standard actor-critic-based algorithms. The authors apply the Lagrangian method and the strong duality property brought by Wasserstein distance (Blanchet & Murthy 2019) to convert the min-max robust Bellman equation into a finite-dimensional optimization problem under an implicit $(s,a)$-rectangular assumption. As they demonstrate the existence of a deterministic Markov optimal policy and value-to-go function for the reformulated expected Bellman-form operator, they design the WRAAC algorithm where two neural networks are constructed to act as the critic and actor: the critic serves as the estimator of the value function, the actor aims to attain the robust optimal policy. WRAAC is implemented as a double-loop algorithm: the inner loop determines the extent of perturbations, and the outer loop optimizes the policy as normal procedures. In the inner loop, the algorithm updates the state $z$ that maximizes the penalized value-to-go function (i.e., by the Wasserstein distance with Lagrangian multiplier) and the Lagrangian multiplier $\lambda^{*}$ by sampling. In essence, the inner loop approximates the procedure (42) by Monte Carlo. With the currently optimal $\lambda^{*}$, the outer loop optimizes the critic and actor by the temporal errors. After the actor is updated, new samples are collected from the environment using the updated policy, and the process repeats. Sharing a similar idea to Neufeld & Sester (2024), the multiplier $\lambda^{*}$ condenses the impact of the worst-case scenario, which is why we argue that the inner loop certifies the extent of perturbations.
6 Beyond Rectangularity: Coupled Uncertainty Modeling